Constants Part A Coulombs law for the magnitude of the force F between two particles with charges Q and Q separated by a distance d is What is the net force exerted by these two charges on a third charge q3 54.5 nC placed between gi and q at 31.135 m? Your answer may be positive or negative, depending on the direction of the force Express your answer numerically in newtons to three significant figures where K-and View Available Hint(s) E0 8.854 x 10 12 C2/(N . m2) is the permittivity of free space Consider two point charges located on the x axis one charge, -13.0 nC, is located at 1.725 m ; the second charge, q2-3 is at the origin (z-0.0000) 1.0 nC Force on q3 - Submit
Answer #1

force on q3 by q1 = kq1q3/r1^2

= 9*10^9*13*10^-9*54.5*10^-9 / (1.725 - 1.135) ^2

= 1.83*10^-5 N ( to the right)

Force on q3 by q2 = kq2q3 / r2^2

= 9*10^9*31*10^-9*54.5*10^-9 / (1.135)^2

= 1.18*10^-5 N ( to the right)

net force on q3 = 10^-5 ( 1.83 + 1.18) = 3.01*10^-5 N (to the right)